OpenStudy (anonymous):

PLEASE CHECK MY ANSWER :) Problem: 1/2x-8=-3/2x+2 Answer: -13.3 If wrong please explain why.

OpenStudy (texaschic101):

1/2x - 8 = -3/2x + 2 -- get rid of fractions by multiplying by common denominator of 2 x - 16 = -3x + 4 x + 3x = 4 + 16 4x = 20 x = 20/4 x = 5

OpenStudy (texaschic101):

Its the way the question is written..... Is it : (1/2x) - 8 = (-3/2x) + 2 or 1/(2x - 8) = -3/(2x + 2)

OpenStudy (anonymous):

The problem is 1/2x-8=-3/2x+2

OpenStudy (anonymous):

so the first one

OpenStudy (texaschic101):

so there is no parenthesis in the problem ??

OpenStudy (anonymous):

no there is not

OpenStudy (texaschic101):

@Loser66 .....

OpenStudy (texaschic101):

with no parenthesis, the problem can be interpreted 2 ways giving you 2 different answers. Lets see what loser says

OpenStudy (texaschic101):

@ganeshie8 ...what would you do

OpenStudy (anonymous):

can I ask you another question while waiting? @texaschic101

ganeshie8 (ganeshie8):

Exatly! as the question stands, the solution looks good to me! Sure, OP needs to know how to use latex/parenthesis properly..

OpenStudy (texaschic101):

thx ganshie.....so x = 5

OpenStudy (texaschic101):

oops...spelled it wrong...meant ganeshie

OpenStudy (anonymous):

so X=5

OpenStudy (texaschic101):

yes

OpenStudy (anonymous):

Thank You so much!!!!!!!!

OpenStudy (texaschic101):

just remember, if you have a problem like this...with fractions, you can get rid of the fractions by multiplying everything by the common denominator

OpenStudy (texaschic101):

or you can just solve it leaving the fractions in....whatever is easiest for you